# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio Choice

### v5

@@ Appendix A Derivation of the Adjoint Diffusion Coefficient via Itô’s Lemma
This derivation provides the explicit form of the adjoint diffusion coefficient $\mathbf{Z}_{t}^{\lambda X}$, which is fundamental for understanding the structure of the optimal portfolio policy (11) and for constructing the Pontryagin adjoint equations.
Appendix B Analytic Solutions for Affine Models
This appendix provides details on the derivation of analytic solutions for the portfolio optimization problems presented in Section 2.3 under specific assumptions (no intermediate consumption, CRRA utility), typically following the Hamilton-Jacobi-Bellman (HJB) approach with an exponential-quadratic value function ansatz. We also show how these solutions connect to the Pontryagin framework.
B.1 Analytic Solution for the Multi-Asset OU Model (Section 2.3.1)
For simplicity, we assume no intermediate consumption, i.e., $C_{t}=0$, and consider an investor who derives utility only from terminal wealth $X_{T}$ under a CRRA objective $U(x)=\frac{x^{1-\gamma}}{1-\gamma}$. Following Kim and Omberg, (1996), the associated Hamilton–Jacobi–Bellman (HJB) equation can be solved by positing a value function of the form:
$V(t,x,Y)=\frac{x^{1-\gamma}}{1-\gamma}\exp\left(g(t)+B(t)Y+\frac{1}{2}C(t)Y^{2}\right).$ (20)
Here, $g(t)$, $B(t)$, and $C(t)$ are deterministic functions of time determined by solving a system of ordinary differential equations (ODEs) derived from the HJB equation, subject to terminal conditions at $t=T$.
From this value function, the costate $\lambda_{t}^{*}=V_{x}$, its derivative with respect to wealth $\partial_{x}\lambda_{t}^{*}=V_{xx}$, and its derivative with respect to the state $\partial_{Y}\lambda_{t}^{*}=V_{xY}$ have the following closed-form expressions on the optimal path $(X_{t}^{*},Y_{t})$:
$\lambda_{t}^{*}(t,X_{t}^{*},Y_{t})$ $=(X_{t}^{*})^{-\gamma}\exp\left(g(t)+B(t)Y_{t}+\frac{1}{2}C(t)Y_{t}^{2}\right),$
$\frac{\partial\lambda_{t}^{*}}{\partial x}(t,X_{t}^{*},Y_{t})$ $=-\gamma(X_{t}^{*})^{-\gamma-1}\exp\left(g(t)+B(t)Y_{t}+\frac{1}{2}C(t)Y_{t}^{2}\right),$
$\frac{\partial\lambda_{t}^{*}}{\partial Y}(t,X_{t}^{*},Y_{t})$ $=(X_{t}^{*})^{-\gamma}\left(B(t)+C(t)Y_{t}\right)\exp\left(g(t)+B(t)Y_{t}+\frac{1}{2}C(t)Y_{t}^{2}\right).$
These expressions lead to remarkably simple ratios that are central to the Pontryagin framework’s characterization of the optimal portfolio:
$\frac{\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma},\qquad\frac{\partial_{Y}\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma}\left(B(t)+C(t)Y_{t}\right).$
Substituting these ratios into the general Pontryagin first-order condition for the optimal portfolio $\bm{\pi}_{t}^{*}$ (which equates the marginal contribution to the Hamiltonian from investing in risky assets to zero, cf. Eq. (10) derived in Section 2.2) confirms the explicit analytic solution derived via the HJB approach:
$\bm{\pi}^{*}(t,Y)=\frac{1}{\gamma}\,\Sigma^{-1}\Bigl{[}\underbrace{(\bm{\alpha}\,\bm{\sigma})\,Y}_{\text{Myopic}}+\underbrace{\sigma_{Y}\,(\bm{\sigma}\,\bm{\rho})\,\bigl{(}B(t)+C(t)\,Y\bigr{)}}_{\text{Hedging}}\Bigr{]},$
where $\Sigma=\mathrm{diag}(\bm{\sigma})\,\Psi\,\mathrm{diag}(\bm{\sigma})$, $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{n})^{\top}$, $\bm{\rho}=(\rho_{1},\dots,\rho_{n})^{\top}$, $(\bm{\alpha}\,\bm{\sigma})$ is the vector $(\alpha_{1}\sigma_{1},\dots,\alpha_{n}\sigma_{n})^{\top}$, and $(\bm{\sigma}\,\bm{\rho})$ is the vector $(\sigma_{1}\rho_{1},\dots,\sigma_{n}\rho_{n})^{\top}$. The derivation confirms that the first term captures the myopic demand driven by the instantaneous risk premium $(\bm{\mu}(Y)-r\mathbf{1})=(\bm{\alpha}\,\bm{\sigma})Y$, while the second term represents the intertemporal hedging demand, proportional to $B(t)+C(t)Y$ which arises from the state-dependency of the investment opportunity set captured by $\partial_{Y}\lambda_{t}^{*}$.
Substituting this $\bm{\pi}^{*}$ back into the HJB equation and matching coefficients in powers of $Y$ yields the following Riccati-type ODEs for $B(t)$ and $C(t)$, along with the ODE for $g(t)$. The system is solved backward from the terminal conditions:
$g(T)=0,\quad B(T)=0,\quad C(T)=0.$ (21)
The ODEs are:
$-\frac{dC(t)}{dt}$ $=-2\,\kappa_{Y}\,C(t)+\sigma_{Y}^{2}\,C(t)^{2}+\frac{1-\gamma}{\gamma}\Bigl{[}\beta_{0}+2\,\sigma_{Y}\,C(t)\,\beta_{1}+\sigma_{Y}^{2}\,C(t)^{2}\,\beta_{2}\Bigr{]},$ (22)
$-\frac{dB(t)}{dt}$ $=-\kappa_{Y}\,B(t)+\sigma_{Y}^{2}\,B(t)\,C(t)-\frac{1-\gamma}{\gamma}\Bigl{[}\sigma_{Y}^{2}\,B(t)\,C(t)\,\beta_{2}+\sigma_{Y}\,B(t)\,\beta_{1}\Bigr{]}+\kappa_{Y}\,\theta_{Y}\,C(t),$ (23)
$\frac{dg(t)}{dt}=(1-\gamma)\,r-\frac{1-\gamma}{2\gamma}\Bigl{[}\beta_{0}+2\sigma_{Y}\,\beta_{1}\,B(t)+\sigma_{Y}^{2}\,\beta_{2}\bigl{(}B(t)^{2}+C(t)\bigr{)}\Bigr{]}+\kappa_{Y}\,\theta_{Y}\,B(t)+\tfrac{1}{2}\,\sigma_{Y}^{2}\,C(t).$ (24)
Here, $r$ is the risk-free rate, $\kappa_{Y}$, $\theta_{Y}$, $\sigma_{Y}$ are the OU parameters, $\gamma$ is the risk aversion coefficient, and $\beta_{0}$, $\beta_{1}$, $\beta_{2}$ are quadratic forms involving the model parameters defined as:
$\beta_{0}:=(\bm{\alpha}\,\bm{\sigma})^{\top}\,\Sigma^{-1}\,(\bm{\alpha}\,\bm{\sigma}),\quad\beta_{1}:=(\bm{\alpha}\,\bm{\sigma})^{\top}\,\Sigma^{-1}\,(\bm{\sigma}\,\bm{\rho}),\quad\beta_{2}:=(\bm{\sigma}\,\bm{\rho})^{\top}\,\Sigma^{-1}\,(\bm{\sigma}\,\bm{\rho}).$
By solving the system of ODEs (22), (23), and (24) backward in time from the terminal conditions (21), one obtains the functions $g(t)$, $B(t)$, and $C(t)$. Substituting these into Eq. (20) gives the complete analytic form of the value function consistent with the underlying HJB/PMP framework.
B.2 Analytic Solution for the Multi-Asset Multi-Factor Model (Section 2.3.2)
[…]
where $\mathbf{B}(t)$ is an $k\times 1$ vector, $\mathbf{C}(t)$ is an $k\times k$ symmetric matrix (note: dimensions based on state vector $\mathbf{Y}\in\mathbb{R}^{k}$), and $g(t)$ is a scalar function, all satisfying ODEs derived from the HJB equation, with terminal conditions $g(T)=0$, $\mathbf{B}(T)=\mathbf{0}$ and $\mathbf{C}(T)=\mathbf{0}$.
Define the relevant matrices and vectors:
$\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n}),\;\bm{\sigma}_{Y}\;(\text{factor diffusion }k\times k),\;\Sigma=\bm{\sigma}\Psi\bm{\sigma}.$
where $\Psi\in\mathbb{R}^{n\times n}$ is the correlation matrix for the $n$ risky assets, and let $\rho\in\mathbb{R}^{n\times k}$ be the cross-correlation matrix between these assets and the $k$-dimensional factor Brownian motion $\mathbf{W}^{Y}_{t}$. Also let $\mathbf{A}=\bigl{(}\bm{\alpha}_{1},\dots,\bm{\alpha}_{n}\bigr{)}^{\top}\in\mathbb{R}^{n\times k}$ collect the factor loadings $\bm{\alpha}_{i}\in\mathbb{R}^{k}$ for each asset $i$.
The costate $\lambda_{t}^{*}=V_{x}$, its derivative with respect to wealth $\partial_{x}\lambda_{t}^{*}=V_{xx}$, and its gradient ($k\times 1$ vector) with respect to the state vector $\partial_{\mathbf{Y}}\lambda_{t}^{*}=V_{x\mathbf{Y}}$ are given by:
$\lambda_{t}^{*}(t,X_{t}^{*},\mathbf{Y}_{t})$ $=(X_{t}^{*})^{-\gamma}\exp\left(g(t)+\mathbf{B}(t)^{\top}\mathbf{Y}_{t}+\tfrac{1}{2}\mathbf{Y}_{t}^{\top}\mathbf{C}(t)\mathbf{Y}_{t}\right),$
$\frac{\partial\lambda_{t}^{*}}{\partial x}(t,X_{t}^{*},\mathbf{Y}_{t})$ $=-\gamma(X_{t}^{*})^{-\gamma-1}\exp\left(g(t)+\mathbf{B}(t)^{\top}\mathbf{Y}_{t}+\tfrac{1}{2}\mathbf{Y}_{t}^{\top}\mathbf{C}(t)\mathbf{Y}_{t}\right),$
$\frac{\partial\lambda_{t}^{*}}{\partial\mathbf{Y}}(t,X_{t}^{*},\mathbf{Y}_{t})$ $=(X_{t}^{*})^{-\gamma}\left(\mathbf{B}(t)+\mathbf{C}(t)\mathbf{Y}_{t}\right)\exp\left(g(t)+\mathbf{B}(t)^{\top}\mathbf{Y}_{t}+\tfrac{1}{2}\mathbf{Y}_{t}^{\top}\mathbf{C}(t)\mathbf{Y}_{t}\right).$
These yield the following scalar and vector ratios, which simplify the PMP optimality conditions:
$\frac{\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma},\qquad\frac{\partial_{\mathbf{Y}}\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma}\left(\mathbf{B}(t)+\mathbf{C}(t)\mathbf{Y}_{t}\right).$
Substituting these into the vector form of the PMP first-order condition for the optimal portfolio $\bm{\pi}_{t}^{*}\in\mathbb{R}^{n}$, (cf. Eq. (11) derived in Section 2.2):
$\bm{\pi}_{t}^{*}=-\frac{1}{X_{t}^{*}\partial_{x}\lambda_{t}^{*}}\Sigma^{-1}\left\{\lambda_{t}^{*}(\bm{\mu}(\mathbf{Y}_{t})-r\mathbf{1})+\bm{\sigma}\rho\bm{\sigma}_{Y}\left[(\partial_{\mathbf{Y}}\lambda_{t}^{*})^{\top}\right]\right\},$
where the risk premium vector is $\bm{\mu}(\mathbf{Y}_{t})-r\mathbf{1}=\bm{\sigma}\mathbf{A}\mathbf{Y}_{t}$, we precisely recover the analytic solution obtained from the HJB framework:
$\bm{\pi}^{*}(t,\mathbf{Y})\;=\;\frac{1}{\gamma}\,\Sigma^{-1}\Bigl{[}\,\underbrace{\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}}_{\text{Myopic component (related to }\bm{\mu}-r\mathbf{1})}\;+\;\underbrace{\bm{\sigma}\,\rho\,\bm{\sigma}_{Y}\Bigl{(}\mathbf{B}(t)\;+\;\mathbf{C}(t)\,\mathbf{Y}\Bigr{)}}_{\text{Hedging component (related to }\partial_{\mathbf{Y}}\lambda_{t}^{*})}\Bigr{]}.$
Here, the first bracketed term corresponds to the myopic component, driven by the instantaneous expected excess returns $\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}$. The second bracketed term incorporates intertemporal hedging demands arising from the desire to hedge against unfavorable changes in the investment opportunity set, as captured by the factor sensitivities $\mathbf{B}(t)+\mathbf{C}(t)\mathbf{Y}$ (proportional to the gradient $\partial_{\mathbf{Y}}\lambda_{t}^{*}$) and the correlations encoded in $\rho$, $\Psi$, and $\Phi$ (implicitly via the ODEs for $\mathbf{B}$ and $\mathbf{C}$).
To determine the time-dependent coefficients $\mathbf{B}(t)$ and $\mathbf{C}(t)$, we solve the backward matrix Riccati ODE system derived from substituting the value function ansatz and optimal portfolio into the HJB equation:
$\frac{d\mathbf{C}(t)}{dt}$ $=\kappa_{Y}^{\top}\mathbf{C}(t)+\mathbf{C}(t)\kappa_{Y}-\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}\mathbf{C}(t)$
$\quad-\frac{1-\gamma}{\gamma}\left[(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})+(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{C}(t)\right.$
$\qquad\qquad\left.+\mathbf{C}(t)(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})+\mathbf{C}(t)(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{C}(t)\right],$ (26)
$\frac{d\mathbf{B}(t)}{dt}$ $=\left[\kappa_{Y}^{\top}-\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}-\frac{1-\gamma}{\gamma}\left(\mathbf{C}(t)(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})+(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})\right)\right]\mathbf{B}(t)$
$\quad-\kappa_{Y}\mathbf{C}(t)\bm{\theta}_{Y}.$ (27)
subject to the terminal conditions $\mathbf{C}(T)=\mathbf{0}$ and $\mathbf{B}(T)=\mathbf{0}$.
To obtain the complete value function (25), the scalar function $g(t)$ must also be determined by solving its own ODE backward from the terminal condition $g(T)=0$. This ODE, derived from the HJB equation by collecting terms independent of $\mathbf{Y}$ and consistent with the corrected single-factor case structure, is given by:
$\frac{dg(t)}{dt}$ $=(1-\gamma)r-\frac{1-\gamma}{2\gamma}\operatorname{Tr}\left((\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})\right)-\frac{1-\gamma}{\gamma}\mathbf{B}(t)^{\top}(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})$
$\quad-\frac{1-\gamma}{2\gamma}\left[\mathbf{B}(t)^{\top}(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma}\Sigma^{-1}\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{B}(t)+\operatorname{Tr}\left((\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma}\Sigma^{-1}\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{C}(t)\right)\right]$
$\quad+\mathbf{B}(t)^{\top}\kappa_{Y}\bm{\theta}_{Y}\quad+\frac{1}{2}\operatorname{Tr}(\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}).$
Here, $\Phi=\mathrm{Cov}(d\mathbf{W}^{Y}_{t})/dt$ is the factor covariance matrix used in the factor dynamics $d\mathbf{Y}_{t}$, and $\operatorname{Tr}(\cdot)$ denotes the trace of a matrix.
Solving these coupled ODEs for $\mathbf{C}(t)$, $\mathbf{B}(t)$, and $g(t)$ backward in time from $T$ to $t$ and substituting the results into Eq. (25) gives the closed-form value function, while substituting $\mathbf{B}(t)$ and $\mathbf{C}(t)$ into $\bm{\pi}^{*}(t,\mathbf{Y})$ gives the optimal strategy, explicitly capturing both myopic and intertemporal hedging components in a multi-asset, multi-factor setting.
Appendix C Complete and Rigorous Proofs for the Policy‑Gap Bound (Theorem 1)
This appendix provides the complete, self‑contained analytic backbone for Theorem 1. All proof sketches have been expanded into detailed arguments, with explicit reliance on parabolic maximal‑regularity and monotone‑operator theory.
C.1 Table of Symbols
@@
$\delta_{\mathrm{BPTT}}$ Costate estimation error (Def. 1)
$\underline{\Lambda}$ Lower eigenvalue from Lemma 4
$C_{S},C_{v}^{\prime\prime}$ PDE stability / coupled‑error constants
$L_{F},C_{\text{tot}}$ Final error amplification constants
C.2 Preliminaries and Notation
Throughout, we set the primary costate
$\lambda:=\nabla_{x}V^{\pi,c}.$
For any admissible control $(\pi,c)$ the sub‑value function $V^{\pi,c}$ solves
$-\partial_{t}V^{\pi,c}=\mathcal{L}_{\pi,c}[V^{\pi,c}]+f_{\pi,c},\qquad V^{\pi,c}(T,\cdot)=g,$
with oblique‑derivative boundary $\partial_{\mathbf{n}}V=0$ on $\partial\Omega$ to match the Skorokhod reflection (Lieberman, , 1996). Error quantities on $Q_{T}:=(0,T)\times\Omega$:
$h$ $:=(a_{\pi,c}-a_{*},\,b_{\pi,c}-b_{*}),$ $\quad E_{h}$ $:=\|h\|_{L^{q,p}},$
$E_{f}$ $:=\bigl{\|}e^{-\delta t}(U(c^{*})-U(c))\bigr{\|}_{L^{q,p}},$ $\quad E_{v}$ $:=\|\nabla V^{\pi,c}-\nabla V^{*}\|_{L^{q,p}},$
$\varepsilon$ $:=\|\mathcal{R}_{\mathrm{FOC}}\|_{L^{q,p}}.$
C.3 Minimal Assumptions
Assumption 6 (Baseline Well‑Posedness).
Fix $T>0$, a $C^{1,1}$ domain $\Omega$, mixed‑norm exponents $p>d+2,\;q\geq 2$ and a compact cylinder
$\mathcal{D}:=[0,T]\times[\underline{x},\bar{x}]\times\{\|y\|\leq R_{Y}\}.$
There exist positive constants $\underline{\nu},\overline{\nu},L_{\mathrm{coef}},C_{\mathrm{MR}},\eta_{0}$ such that

### v6

@@ Appendix B Derivation of the Adjoint Diffusion Coefficient via Itô’s Lemma
This derivation provides the explicit form of the adjoint diffusion coefficient $\mathbf{Z}_{t}^{\lambda X}$, which is fundamental for understanding the structure of the optimal portfolio policy (3.2) and for constructing the Pontryagin adjoint equations.
Appendix C Proofs
C.1 Proof of Lemma 1
The second term becomes $\lambda_{k+1}^{(i)}(\partial X_{k+1}^{(i)}/\partial X_{k}^{(i)})$. Differentiating the discrete wealth dynamics, which includes the consumption term $C_{k}^{(i)}\Delta t$, with respect to $X_{k}^{(i)}$ yields:
Substituting these components and rearranging the terms for $\lambda_{k}^{(i)}$ yields the expression in Eq. (1).
C.2 Proof of Lemma 2
The product inside the expectation is an inner product of vectors. The term involving the adapted part, $\mathbb{E}[\lambda_{k+1}(\dots)|\mathcal{F}_{t_{k}}]$, is zero as it is the expectation of a martingale increment. The non-zero terms arise from the quadratic covariation of the raw martingale components of $\lambda_{k+1}^{(i)}$ with the wealth shock. Using the standard properties $\mathbb{E}[(\Delta\mathbf{W}_{k}^{X})(\Delta\mathbf{W}_{k}^{X})^{\top}|\mathcal{F}_{t_{k}}]=I\Delta t$ and $\mathbb{E}[(\Delta\mathbf{W}_{k}^{Y})(\Delta\mathbf{W}_{k}^{X})^{\top}|\mathcal{F}_{t_{k}}]=\bm{\rho}_{Y\!X}\Delta t$, where $\bm{\rho}_{Y\!X}$ is the cross-correlation matrix, yields the stated result.
C.3 Proof of Theorem 3
We begin by taking the conditional expectation $\mathbb{E}[\cdot\mid\mathcal{F}_{t_{k}}]$ of the entire pathwise recurrence in Eq. (1) from Lemma 1. The left-hand side becomes $\lambda_{k}$ by definition. On the right-hand side, the expectations of the drift terms are straightforward. The crucial step is handling the stochastic term $\mathbb{E}[\lambda_{k+1}^{(i)}(\dots)|\mathcal{F}_{t_{k}}]$, for which we substitute the result from Lemma 2. This transforms the stochastic term into the additional drift component involving the effective process $\mathbf{Z}^{\lambda X}_{k}$. After collecting all drift terms and separating the martingale components, we arrive at the discrete-time BSDE presented in Eq. (3). The final step involves a direct algebraic comparison of this drift term with the definition of the Hamiltonian $\mathcal{H}$ from Section 3.2, which confirms their identity.
C.4 Proof of Theorem 4.2
This appendix provides a complete, self-contained analytic backbone for Theorem 4.2. All proof sketches are expanded into detailed arguments, relying on parabolic maximal regularity (MR) with VMO (vanishing mean oscillation) coefficients and the strong monotonicity of the PMP first-order condition (FOC) map. The reflected SDE framework is used purely as a localization device; passing to expanding domains recovers the original non-reflected problem (Karatzas and Shreve, , 1991; Lieberman, , 1996).
C.4.1 Table of Symbols
@@
$\delta_{\mathrm{BPTT}}$ Costate estimation error (Def. 2)
$\underline{\kappa}$ Strong monotonicity modulus of the FOC map (Lemma 4)
$C_{S},C_{v}^{\prime\prime}$ PDE stability / coupled-error constants
$L_{G},L_{F},C_{\text{tot}}$ FOC-map Lipschitz / final amplification constants
$\tau_{\star}$ Contraction slab length (distinct from numerical step $\Delta t$)
C.4.2 Preliminaries and Notation
Throughout, set the primary costate $\lambda:=\partial_{x}V^{\pi,c}$. For any admissible control $(\pi,c)$ the sub-value function $V^{\pi,c}$ solves
$\partial_{t}V^{\pi,c}\;=\;\mathcal{L}_{\pi,c}[V^{\pi,c}]+f_{\pi,c},\qquad V^{\pi,c}(T,\cdot)=g,$
on $Q_{T}:=(0,T)\times\Omega$, with oblique-derivative boundary $\partial_{\mathbf{n}}V=0$ on $\partial\Omega$ matching a Skorokhod reflection used for localization; solutions on expanding domains converge to the non-reflected solution (Karatzas and Shreve, , 1991; Lieberman, , 1996).
Define the error quantities on $Q_{T}$ by
$h$ $=(a_{\pi,c}-a_{\ast},\,b_{\pi,c}-b_{\ast}),$ $\quad E_{h}$ $=\|h\|_{L^{q,p}},$
$E_{f}$ $=\bigl{\|}e^{-\delta t}\bigl{(}U(c^{\ast})-U(c)\bigr{)}\bigr{\|}_{L^{q,p}},$ $\quad E_{v}$ $=\|\nabla V^{\pi,c}-\nabla V^{\ast}\|_{L^{q,p}},$
$\varepsilon$ $=\|\mathcal{R}_{\mathrm{FOC}}\|_{L^{q,p}},$
Convention on $Z_{\lambda X}$.
We use
$Z_{\lambda X}\;=\;(\partial_{x}\lambda)\,X\,\pi^{\top}\sigma\;+\;(\partial_{Y}\lambda)\,\sigma_{Y}\,\rho^{\top},\qquad\sigma\,Z_{\lambda X}^{\top}\;=\;X\bigl{(}-\partial_{x}\lambda\bigr{)}\,\Sigma\,\pi\;+\;\sigma\rho\sigma_{Y}^{\top}(\partial_{Y}\lambda)^{\top},$
consistent with the portfolio FOC (cf. main text Eq. (11)); see Yong and Zhou, (2012) for PMP–HJB links.
C.4.3 Minimal Assumptions
Assumption 5 (Baseline well-posedness and regularity).
Fix $T>0$, a bounded $C^{1,1}$ domain $\Omega$, mixed-norm exponents $p>d+2$ and $q\geq 2$, and a compact cylinder $\mathcal{D}=[0,T]\times[\underline{x},\bar{x}]\times\{\|y\|\leq R_{Y}\}$. There exist constants $\underline{\nu},\overline{\nu},L_{\mathrm{coef}},C_{\mathrm{MR}},\eta_{0}>0$ such that: